Levine Inc., which produces a single product, has prepared the following standard cost sheet for one unit of the product.

Direct materials (7 pounds at $1.60 per pound) $11.20
Direct labor (1 hours at $10.00 per hour) $10.00


During the month of April, the company manufactures 350 units and incurs the following actual costs.

Direct materials purchased and used (2,000 pounds) $3,600
Direct labor (370 hours) $3,663


Compute the total, price, and quantity variances for materials and labor.

Total materials variance $enter a dollar amount select a type of the variance FavorableUnfavorableNeither favorable nor unfavorable
Materials price variance $enter a dollar amount select a type of the variance FavorableNeither favorable nor unfavorableUnfavorable
Materials quantity variance $enter a dollar amount select a type of the variance UnfavorableNeither favorable nor unfavorableFavorable
Total labor variance $enter a dollar amount select a type of the variance FavorableNeither favorable nor unfavorableUnfavorable
Labor price variance $enter a dollar amount select a type of the variance FavorableUnfavorableNeither favorable nor unfavorable
Labor quantity variance $enter a dollar amount select a type of the variance FavorableNeither favorable nor unfavorableUnfavorable
Answer #1

Total materials variance = (11.20*350) - 3600

= 320 Favourable

Materials Price Variance = (SP-AP)*AQ = (1.60-3600/2000)*2000

= 400 Unfavorable

Materials Quantity Variance = (SQ-AQ) *SP = (350*7 - 2000)*1.60

= 720 Favourable

Total labor variance = (10*350) - 3663

= 163 Unfavorable

Labor Rate Variance = (SR-AR) *AH = (10-3663/370)*370

= 37 Favourable

Labor quantity Variance = (SH-AH) *SR = (350*1 - 370)*10

= 200 Unfavorable